The mass ofa nonvolatile solute of malar mass 40g/mol that should be dissolved in114g of octane to lower its vapour press.by 20% is.........
 
a) 10 g                      b)11.4 g
c)9.8 g                      d)12.8 g

Grade:12th pass

1 Answers

Vikas TU
7 years ago
From partial pressure for non volatile solute formulae:
(P – xP)/P = nA/nB
(P – 0.2P)/P = (w/40)/(114/114)
0.8 = w/40
w = 32 g.